# Advances in Robust Federated Learning: Heterogeneity Considerations

### v1

@@ III-C3 Hierarchical Federation Learning
Liu et al. [135] propose a client-edge cloud hierarchical joint learning system, HierFAVG, an algorithm that allows multiple edge servers to perform partial model aggregation, which allows for better communication and utility tradeoffs. Abad et al. [10] collaboratively learn global models by sharing local updates of model parameters instead of their dataset, and in this way use a hierarchical joint learning scheme to significantly reduce communication latency without sacrificing accuracy. Briggs et al. [136] separated clusters of clients by introducing a hierarchical clustering step that uses the similarity of the client’s local updates to the global joint model. Xu et al. [137] consider a hierarchical FL system and propose a joint problem of edge aggregation interval control and resource allocation to minimize a weighted sum of training loss and training delay. Lim et al. [138] use an evolutionary game to model workers’ edge association strategies and propose a hierarchical game framework to study the dynamics of edge association and resource allocation in self-organizing HFL networks. Abdellatif et al. [52] investigated the potential of hierarchical FL in IoT heterogeneous systems and proposed a user assignment and resource allocation method on a hierarchical FL architecture for IoT heterogeneous systems. Zhou et al. [139] proposed a unique clustering-based approach for participant selection using social context data. By creating different groups of edge participants and performing group-specific joint learning, the models of various edge groups are further aggregated to enhance the robustness of the global model. Ma et al. [140] proposed FedUC, which clusters workers through a hierarchical aggregation structure to solve the edge heterogeneity problem and reduce the communication consumption of FL. Deng et al. [141] proposed a communication cost minimization problem in the HFL framework by making decisions on edge enhancer selection and node edge association in order to achieve the target accuracy of the total communication cost required for model learning.
@@ III-C4 Decentralized Federated Learning
Decentralized federated learning (DFL) is an approach that does not rely on a central server to manage model training. Compared with traditional centralized federated learning, decentralized methods provide higher privacy protection.
@@
Li et al. [142] developed a decentralized FL framework BFLC based on blockchain with committee consensus. It uses blockchain for global model storage and local model update exchange [143], and reduces malicious client attacks through committee verification composed of honest clients. Lim et al. [144] proposed a deep learning-based auction mechanism to derive the valuation of the services of each cluster leader. ye et al. [145] proposed a robust decentralized stochastic gradient descent federated method, called soft DSGD, to solve the problem of unreliable transmission process. Pappas et al. [146] designed a fully decentralized FL framework based in part on the Inter Planetary File System (IPFS). By using IPLS and connecting to the corresponding private IPFS network, any party can initiate the training process of a machine learning model, or join an ongoing training process initiated by the other party. Liu et al. [147] proposed a general DFL framework that regularly implements multiple local updates and multiple inter-node communications to strike a balance between communication efficiency and model consistency. Kalra et al. [148] proposed an efficient communication scheme for decentralized FL, called ProxyFL. Each participant in ProxyFL maintains two models, one is a private model and the other is a public shared proxy model. Designed to protect participant privacy. The agent model allows efficient information exchange between participants without the need for a centralized server. HegedHus et al. [149] believe that Gossip learning is a decentralized alternative to FL, so they designed a DFL training algorithm based on the Gossip protocol. Qu et al. [150] proposed a new architecture called DFL for drone networks (DFL-UN), which implements FL in drone networks without the need for a central entity. Beltran et al. [151] introduced a novel platform, Fedstar, which allows users to create federations by customizing parameters. Fedstar aims to train FL models in different federations of physical or virtual devices in a decentralized, semi-decentralized and centralized manner. Zhou et al. [152] proposed a decentralized federated trusted average (DeFTA), as a plug-and-play replacement for FedAvg, brings immediate improvements to security, scalability, and fault tolerance in the FL process. Cheng et al. [153] used a decentralized and trusted aggregation strategy to introduce DeTA, a decentralized FL system architecture, and designed a defense-in-depth model.
@@ IV Privacy Protection in Heterogeneous Federated Learning
In this section, we will focus on the privacy and security issues in heterogeneous FL. As is well recognized, FL frameworks are deployed over the internet. The complex and unpredictable network environment poses potential vulnerabilities for attacks on the FL process. This underscores the necessity of employing protective measures within FL applications. Since the inception of the concept of FL, researchers have published a range of attack methods targeting the exposure of certain stages to the network environment. These methods can launch attacks at various stages, such as dataset preparation [154], model transmission [155], and model aggregation [156], with malicious intentions of privacy theft [157] or interference [158] with the training process. Among them, model inversion attacks [12], which involve eavesdropping on the model transmission channel to intercept the model and attempt to reconstruct the client’s private data from the captured information, pose significant threats to FL frameworks. In some cases, it is even possible to fully reconstruct the private dataset from the model. For attack vectors such as model inversion attacks, model protection is particularly critical. Through years of research, model protection can be categorized into three types [15]: differential privacy, secure multi-party computation, and model tracing, each providing pre-emptive or post-emptive protection for the transmitted models from different perspectives. It is noteworthy that different types of protective methods will have varying impacts on the trade-off [159, 160] between privacy, utility, and efficiency, thus analyzing privacy-preserving strategies is crucial for privacy-utility trade-offs in heterogeneous FL. We show common FL privacy-preserving strategies in Fig. 4.
IV-A Differential Privacy
@@
IV-A1 DP and Noising Mechanism
Based on relevant derivations, we generally use a noise mechanism to perturb the transmitted model, thereby making the entire FL mechanism satisfy DP. This mechanism is also known as model perturbation. Before introducing the noise mechanism in detail, we first give the formal definition of DP.
Theorem 1 ( $(\varepsilon,\delta)$-DP).
A randomized mechanism $\mathcal{M}:\cal{X}\mapsto\cal{R}$ with domain $\cal{X}$ and range $\cal{R}$ satisfies $(\varepsilon,\delta)$-DP if, for any two adjacent inputs $x,x^{\prime}\in\cal{X}$ and for any subsets of outputs $\cal{S}\subseteq\cal{R}$ it holds that
$\Pr[\mathcal{M}(x)\in\mathcal{S}]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{M}(x^{\prime})\in\mathcal{S}]+\delta$ (3)
@@ IV-A1 DP and Noising Mechanism
In the context of model perturbation, the model information prepared for transmission undergoes two stages of processing. The first stage is model clipping, which involves setting a threshold $C$. If the norm of the model exceeds this threshold, the model is scaled proportionally until its norm equals the threshold. The second stage involves noising, specifically generating Gaussian noise with a chosen standard deviation $\sigma$ to perturb the model.
@@
The relationship between the noise amplitude $\sigma$ and the privacy budget $\varepsilon$ can be concluded from relevant DP theories, such as Parallel Composition Theorem, Advanced Composition Theorem [162], and so on. Compared to the theories mentioned above, the Moment Accountant [163] theory provides a smaller upper bound for the privacy budget, meaning that this theory can derive a smaller $\varepsilon$ from the same $\sigma$.
There are numerous federated learning frameworks that apply model perturbation. For instance, DP-FedAvg [164], which was initially proposed, first adopted user-level DP to protect client data. NbAFL[165] considered the noise magnitude required for differential privacy protection of model information in both upload and download links. The work of Seif et al. [166] considered the environment of wireless communication and proposed how to implement differential privacy protection in this environment by utilizing the characteristics of wireless communication. CLDP-SGD [167], while maintaining DP protection, considered communication overhead and is a federated learning architecture that applies DP under communication-limited conditions. In addition, some works like Fed-CDP [168] set the object of protection as the training set data, that is, noise is added during local training.
IV-A2 Variant DP
In traditional DP, the relationship between the magnitude of noise and the privacy budget is intricate and not easily described intuitively. To address this issue, Ilya have slightly relaxed the requirements of DP, resulting in the formulation of Rényi Differential Privacy (RDP) [169].
Theorem 2 ($(\alpha,\rho)$-RDP).
Given that $\alpha\in(1,\infty)$ and $\rho>0$, a randomized mechanism $\mathcal{M}$ with domain $\cal{X}$ and range $\cal{R}$ satisfies $(\alpha,\rho)$-RDP if, for any two adjacent inputs $x,x^{\prime}\in\cal{X}$ the $\alpha$-Rényi divergence satisfies
$D_{\alpha}[\mathcal{M}(x)\parallel\mathcal{M}(x^{\prime})]\leftarrow\frac{1}{\alpha-1}\log\mathbb{E}\left[\left(\frac{\mathcal{M}(x)}{\mathcal{M}(x^{\prime})}\right)^{\alpha}\right]\leq\rho$ (5)
where the expectation is taken over $\mathcal{M}(x^{\prime})$.
RDP offers several advantages for federated learning, which are encapsulated in the following points [169]:
  • It is compatible with the Gaussian mechanism, and it can be proven that when employing a Gaussian noise mechanism, the relationship is given by $\sigma^{2}=\frac{(\Delta f)^{2}\alpha}{2\rho}$, where $\Delta f$ refers to the sensitivity of function $f$ and defined as $\Delta f\leftarrow\max_{x,x^{\prime}\,\text{adjacent}}\left\|f(x)-f(x^{\prime})\right\|_{2}$;
  • RDP is amenable to composition, meaning that the outcomes processed sequentially through $(\alpha,\rho_{1})$-RDP and $(\alpha,\rho_{2})$-RDP mechanisms satisfy the composition property of $(\alpha,\rho_{1}+\rho_{2})$-RDP.
  • RDP can be transformed into the classical DP, facilitating comparative analysis of privacy with other mechanisms.
More recent research in the field of federated learning has replaced classical DP with RDP as the basis for privacy. DP-FedSAM [170] is proposed under the consideration of RDP on how to improve the optimizer to enhance model utility. Fed-SMP [171] considers how to apply RDP’s privacy guarantee to the model after sparsification.
IV-A3 DP with Heterogeneity
@@
  • Different clients have varying requirements for privacy, making it difficult to determine the privacy of the entire system;
  • If a client’s privacy budget is too small, the model information from that client is likely to affect model performance [164];
@@
In response to these issues, the academic community has proposed some solutions. PFA [173] classifies clients into public and private categories based on their privacy budgets. After the models of the two types of clients are aggregated separately according to the privacy budget, the model of the private client is projected into the space of the public client’s model. This takes into account the different credibility of information with different privacy levels and does not waste data with stronger privacy. In some cases, the network environment where the federated learning algorithm is deployed also has heterogeneity, and FedSeC [174] has proposed improvements from this perspective. There are also works like DP-SCAFFOLD [175], which advance from the perspective of reducing model utility loss.
IV-B Secure Multi-party Computation
@@
IV-B1 Homomorphic Encryption
Homomorphic encryption (HE) [179] is a specialized cryptographic scheme that imposes additional requirements on certain computational properties to be preserved beyond traditional encryption/decryption functionalities. Different types of HE satisfy distinct properties as follows [180]. Assuming that +,+$\cdot$+ , ⋅ are addition and multiplication on plaintext space and $\oplus$,$\odot$ , ⊙ are ones on cipher space, and $\mathsf{E}$ refers to encryption operation.
  • Addictive Homomorphism. $\mathsf{E}(a+b)=\mathsf{E}(a)\oplus\mathsf{E}(b)$ for any plaintexts $a,b$;
@@ IV-B1 Homomorphic Encryption
A cryptographic scheme satisfying only one property is termed partially HE, whereas one that satisfies both properties is named fully HE. When HE is employed in federated learning [181], the model can be encrypted before transmission, preventing even observers or servers from accessing model information. Building upon this foundation, leveraging the properties of homomorphic encryption, servers can perform aggregation operations on the model without acquiring the model information. In the current landscape of common homomorphic encryption schemes, the Paillier cryptosystem [182], as an additive homomorphic encryption scheme, was introduced relatively early and is extensively used in federated learning. The GSW cryptosystem [183] was the first to propose bootstrap, enabling homomorphic multiplication to transition from a finite number of times to an infinite number, achieving full homomorphic encryption in the truest sense. The CKKS cryptosystem [184], built upon the GSW cryptosystem, enables direct encryption of real number vectors, making it more applicable to machine learning scenarios.
However, different clients may hold different keys, and in Homomorphic Encryption (HE), the same key must be used for encryption to perform ciphertext computation. This poses a challenge for HE in heterogeneous scenarios in federated learning. An effective solution is key distribution, where all client keys are managed and distributed by a trustworthy third party (TTP). This scheme is easy to implement, but the security of the system relies on whether the TTP is truly trustworthy. Another method is called key agreement, which does not require a TTP. Each client communicates on their own, negotiating with the keys held by themselves and the other party to generate a consistent session key. Currently, there is a framework called xMK-CKKS [185], designed based on the key agreement scheme, that can merge all client public keys into one key for CKKS encryption.
IV-B2 Functional Encryption

### v2

@@ III-C3 Hierarchical Federated Learning
Liu et al. [149] propose a client-edge cloud hierarchical FL system, HierFAVG, an algorithm that allows multiple edge servers to perform partial model aggregation, which allows for better communication and utility tradeoffs. Abad et al. [11] collaboratively learn global models by sharing local updates of model parameters instead of their dataset, and in this way use a hierarchical joint learning scheme to significantly reduce communication latency without sacrificing accuracy. Briggs et al. [150] separated clusters of clients by introducing a hierarchical clustering step that uses the similarity of the client’s local updates to the global joint model. Xu et al. [151] consider a hierarchical FL system and propose a joint problem of edge aggregation interval control and resource allocation to minimize a weighted sum of training loss and training delay. Lim et al. [152] use an evolutionary game to model workers’ edge association strategies and propose a hierarchical game framework to study the dynamics of edge association and resource allocation in self-organizing HFL networks. Abdellatif et al. [95] investigated the potential of hierarchical FL in IoT heterogeneous systems and proposed a user assignment and resource allocation method on a hierarchical FL architecture for IoT heterogeneous systems. Zhou et al. [153] proposed a unique clustering-based approach for participant selection using social context data. By creating different groups of edge participants and performing group-specific joint learning, the models of various edge groups are further aggregated to enhance the robustness of the global model. Ma et al. [154] proposed FedUC, which clusters workers through a hierarchical aggregation structure to solve the edge heterogeneity problem and reduce the communication consumption of FL. Deng et al. [60] proposed a communication cost minimization problem in the HFL framework by making decisions on edge enhancer selection and node edge association in order to achieve the target accuracy of the total communication cost required for model learning.
@@ III-C4 Decentralized Federated Learning
Decentralized federated learning (DFL) is an approach that does not rely on a central server to manage model training. Compared with traditional centralized FL, decentralized methods provide higher privacy protection.
@@
Li et al. [78] developed a decentralized FL framework BFLC based on blockchain with committee consensus. It uses blockchain for global model storage and local model update exchange [155], and reduces malicious client attacks through committee verification composed of honest clients. Lim et al. [156] proposed a deep learning-based auction mechanism to derive the valuation of the services of each cluster leader. Ye et al. [61] proposed a robust decentralized stochastic gradient descent federated method, called soft DSGD, to solve the problem of unreliable transmission process. Pappas et al. [157] designed a fully decentralized FL framework based in part on the Inter Planetary File System (IPFS). By using IPLS and connecting to the corresponding private IPFS network, any party can initiate the training process of a ML model, or join an ongoing training process initiated by the other party. Liu et al. [158] proposed a general DFL framework that regularly implements multiple local updates and multiple inter-node communications to strike a balance between communication efficiency and model consistency. Kalra et al. [159] proposed an efficient communication scheme for decentralized FL, called ProxyFL. Each participant in ProxyFL maintains two models, one is a private model and the other is a public shared proxy model. Designed to protect participant privacy. The agent model allows efficient information exchange between participants without the need for a centralized server. HegedHus et al. [160] believe that Gossip learning is a decentralized alternative to FL, so they designed a DFL training algorithm based on the Gossip protocol. Qu et al. [161] proposed a new architecture called DFL for drone networks (DFL-UN), which implements FL in drone networks without the need for a central entity. Beltran et al. [162] introduced a novel platform, Fedstar, which allows users to create federations by customizing parameters. Fedstar aims to train FL models in different federations of physical or virtual devices in a decentralized, semi-decentralized and centralized manner. Zhou et al. [163] proposed a decentralized federated trusted average (DeFTA), as a plug-and-play replacement for FedAvg, brings immediate improvements to security, scalability, and fault tolerance in the FL process. Cheng et al. [164] used a decentralized and trusted aggregation strategy to introduce DeTA, a decentralized FL system architecture, and designed a defense-in-depth model.
III-C5 Asynchronous Federated Learning
Traditional FL typically employs a synchronous aggregation strategy. When devices are heterogeneous, resource utilization is constrained because the aggregation process in each training round must wait for the slower devices. Asynchronous FL [66] is a distributed ML approach designed to address the inefficiencies in training caused by disparities in computational power and communication latency among different participants in traditional FL. In asynchronous FL, participants are not required to synchronize their model updates [165]. Instead, they can independently conduct local training and update their models based on their computational speed and network conditions. The central server continuously receives updates from various participants and integrates these updates into the global model.
Chen et al. [85] allow each client to run the stochastic gradient algorithm without coordinating with other clients, and thus is suitable for intermittent connections of clients. The method further uses a new perturbed local embedding technique to ensure data privacy and improve communication efficiency. Sprague et al. [166] proposed a new asynchronous FL algorithm and investigated the speed of convergence when distributed over multiple edge devices under hard data constraints, and applied asynchronous FL to challenging geospatial applications. Nguyen et al. [53] address the inability of synchronous FL to efficiently scale to hundreds of parallel client trainings by proposing a new buffered asynchronous aggregation method, FedBuff, which is independent of the choice of optimizer and combines the best features of both synchronous and asynchronous FL. Chen et al. [167] proposed an asynchronous online FL (ASO Fed) framework in which edge devices use continuously streaming local data for online learning and a central server aggregates model parameters from clients. Wang et al. [168] propose a new asynchronous FL framework that designs a new centralized fusion algorithm for determining the fusion weights during global updates by taking into account the possible delays in training and uploading local models. To solve the problems of drastically reduced training accuracy and slow convergence in asynchronous FL, Zhang et al. [169] proposed TimelyFL, a heterogeneous perceptual asynchronous FL framework with adaptive partial training.
@@ IV Privacy Protection in Heterogeneous Federated Learning
In this section, we will focus on the privacy and security issues in heterogeneous FL. As is well recognized, FL frameworks are deployed over the internet. The complex and unpredictable network environment poses potential vulnerabilities for attacks on the FL process. This underscores the necessity of employing protective measures within FL applications. To address the security and privacy issuses, protective methods are particularly critical. Through years of research, model protection can be categorized into three types [16]: differential privacy, secure multi-party computation, and model tracing, each providing pre-emptive or post-emptive protection for the transmitted models from different perspectives. We show common FL privacy-preserving strategies in Fig. 6.
IV-A Potential Threats
Any stage of FL exposed in a network environment is susceptible to security and privacy threats. We can categorize these threats based on the methods and objectives of the attacks as follows.
IV-A1 Byzantine Attacks
With such attacks, malicious nodes exhibit arbitrary behavior during the training process to achieve certain malicious objectives. The following attack methods are most representitive.
  • Gradient Poisoning Attacks [170]: The attacker manipulates the training direction by uploading carefully crafted malicious gradients.
  • Gradient Reversal Attacks [171]: The attacker reverses the local gradients before uploading, causing the model to learn in the wrong direction.
  • Replaying Attacks [172]: The attacker captures and replays legitimate communication messages (such as local gradients) to disrupt the learning process.
  • Sybil Attacks [173]: The attacker creates a large number of fake nodes to hijack the consensus in FL, amplifying the impact of the aforementioned attacks.
Furthermore, the aggregator in FL could also be a Byzantine node, with the capability to directly manipulate the training process. In such cases, stronger security measures and robust authorization methods are required.
IV-A2 Interference Attacks
Adversaries employing such attacks do not interfere with the training process but instead eavesdrop on the information transmitted between nodes, attempting to infer the embedded private information. There are also several well-known inference attacks:
  • Membership Inference Attacks [15]: An external eavesdropper intercepts the information transmitted between nodes and attempts to infer embedded private details.
  • Attribute Inference Attacks [174]: The attacker aims to infer certain attributes embedded in the training data, even if these attributes were not explicitly used in the training process. By analyzing the model’s output or behavior, the attacker can infer specific attributes.
  • Model Inversion Attacks [13]: The attacker attempts to reconstruct the client’s private data from the captured information, and it is even possible to fully reconstructbthe private dataset from the model with the assistant of GAN.
IV-A3 Challanges from Heterogeneity
The attacks mentioned before challenge the robustness of FL systems. Even worse, the heterogenous scenarios make it more difficult to detect or defend against such attacks. Data heterogeneity causes the privacy information of individual clients more susceptible to inteference attacks and that malicious Byzantine gradients are more challenging to identify [175]. Communication and device heterogeneity indicate that some communication links in FL systems are more vulnerable to attacks, primarily depending on the protective strategies adopted by the devices. To solve these issues, researchers propose place their interest on heterogeneity-independent protective methods, which is able to be flexibly configured as plugins within FL systems, such as the DP and SMC techniques discussed later.
IV-A4 Compromise on Performence
Even though protection methods can mitigate the adverse effects of the aforementioned attacks, the associated costs persist. Model perturbation-based methods enable relatively quick protection but reduce model accuracy, whereas cryptographic approaches maintain model performance nearly unaffected [67]. However, their computational and communication overhead remains prohibitively high in distributed environments [176]. Existing theory has demonstrated that in FL, it is impossible to simultaneously optimize privacy, utility, and efficiency; a trade-off is inevitable [177]. Therefore, the choice of protection methods introduced later should be guided by the preferences of administrators or users in FL systems.
IV-B Differential Privacy
@@
IV-B1 Definitions
Before introducing the noise mechanism in detail, we first give the formal definition of DP.
Definition 1 ( $(\varepsilon,\delta)$-DP).
A randomized mechanism $\mathcal{M}$ with range $\cal{R}$ satisfies $(\varepsilon,\delta)$-DP if, for any two adjacent datasets $\cal{D},\cal{D^{\prime}}$ and for any subsets of outputs $\cal{S}\subseteq\cal{R}$ it holds that
$\Pr[\mathcal{M}(\cal{D})\in\mathcal{S}]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{M}(\cal{D^{\prime}})\in\mathcal{S}]+\delta$ (3)
@@ IV-B1 Definitions
Definition 1 represents the privacy of a FL algorithm by conducting the probability of an attacker interferencing successfully whether a member is in the dataset during training. In the context of FL, adjacent datasets is ones that differ only in the data held by a single client, where such DP is called user-level.
IV-B2 Model Perturbation
Based on relevant derivations, we generally use a noise mechanism to perturb the transmitted model, thereby making the entire FL mechanism satisfy DP. This mechanism is also known as model perturbation, which has two stages. The first stage is model clipping, which involves setting a threshold $C$. If the norm of the model exceeds this threshold, the model is scaled proportionally until its norm equals the threshold. The second stage involves noising, specifically generating Gaussian noise with a chosen standard deviation $\sigma$ to perturb the model.
@@ IV-B2 Model Perturbation
The relationship between the noise amplitude $\sigma$ and the privacy budget $\varepsilon$ can be concluded from relevant DP theories, such as Advanced Composition Theorem [179], Moment Accountant [180] and variant DPs like RDP [181] or zCDP [182]. All the theories mentioned before persue a smaller upper bound for the privacy budget, meaning that they can derive a smaller $\varepsilon$ from the same noise amplitude.
There are numerous federated learning frameworks that apply model perturbation. For instance, DP-FedAvg [67], which was initially proposed, first adopted user-level DP to protect client data. NbAFL [68] considered the noise magnitude required for differential privacy protection of model information in both upload and download links. Seif et al. [69] consider the environment of wireless communication and propose how to implement differential privacy protection in this environment by utilizing the characteristics of wireless communication.
IV-B3 DP with Heterogeneity
@@
  • Different clients have varying requirements for privacy, making it difficult to determine the privacy of the entire system. Such phenomenon can be called privacy heterogeneity.
@@
In response to these issues, the academic community has proposed some solutions. PFA [70] classifies clients into public and private categories based on their privacy budgets, and then utilize the different roles for better accuracy. In some cases, the network environment where the federated learning algorithm is deployed also has heterogeneity, and FedSeC [184] has proposed improvements from this perspective. There are also works like DP-SCAFFOLD [71], which advance from the perspective of reducing model utility loss.
IV-C Secure Multi-party Computation
@@
IV-C1 Homomorphic Encryption
Homomorphic encryption (HE) [188] is a specialized cryptographic scheme that imposes additional requirements on certain computational properties to be preserved beyond traditional encryption/decryption functionalities. Assuming that +,+$\cdot$+ , ⋅ are addition and multiplication on plaintext space and $\oplus$,$\odot$ , ⊙ are ones on cipher space, and $\mathsf{E}$ refers to encryption operation.
  • Additive Homomorphism. $\mathsf{E}(a+b)=\mathsf{E}(a)\oplus\mathsf{E}(b)$ for any plaintexts $a,b$;
@@ IV-C1 Homomorphic Encryption
An HE scheme satisfies at least one of the properties. When HE is employed in federated learning [72], the model can be encrypted before transmission, preventing even observers or servers from accessing model information. Building upon this foundation, leveraging the properties of homomorphic encryption, servers can perform aggregation operations on the model without acquiring the model information.
In the current HE, there are two schemes are useful, Paillier cryptosystem [189], and CKKS cryptosystem [190]. The former encrypts integers and remain additional homomorphism and the latter enables direct encryption of real number vectors, and perform homomorphic addition and multiplication, making it more applicable to machine learning scenarios.
IV-C2 Functional Encryption